OpenStudy (anonymous):

how do you get 1/4 of 140?

jimthompson5910 (jim_thompson5910):

think of 140 as 140/1

jimthompson5910 (jim_thompson5910):

so what we have is \[\Large \frac{1}{4} \times \frac{140}{1}\]

jimthompson5910 (jim_thompson5910):

multiply the fractions by multiplying straight across (numerator times numerator, denominator times denominator) be sure to reduce as much as possible